### v8

# Solving the Clustering Reasoning Problems by Modeling a Deep-Learning-Based Probabilistic Model

Abstract
Visual abstract reasoning problems pose significant challenges to the perception and cognition abilities of artificial intelligence algorithms, demanding deeper pattern recognition and inductive reasoning beyond mere identification of explicit image features. Research advancements in this field often provide insights and technical support for other similar domains. In this study, we introduce PMoC, a deep-learning-based probabilistic model, achieving high reasoning accuracy in the Bongard-Logo, which stands as one of the most challenging clustering reasoning tasks. PMoC is a novel approach for constructing probabilistic models based on deep learning, which is distinctly different from previous techniques. PMoC revitalizes the probabilistic approach, which has been relatively weak in visual abstract reasoning. As a bonus, we also designed Pose-Transformer for complex visual abstract reasoning tasks. Inspired by capsule networks, it focuses on positional relationships in image data, boosting accuracy when combined with PMoC. Our Pose-Transformer effectively addresses reasoning difficulties associated with changes in the position of entities, outperforming previous models on RAVEN dataset, and the PGM dataset. RAVEN and PGM represent two significant progressive pattern reasoning problems. Finally, considering the deployment difficulties of Pose-Transformer, we introduced Straw-Pose-Transformer, a lightweight version. This study contributes to enhancing the capabilities of artificial intelligence in abstract reasoning, cognitive pattern, and probabilistic modeling of complex systems.
Index Terms:
Visual abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Deep-Learning-Based Probabilistic Model.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Deep neural networks have achieved remarkable success in various fields, including computer vision [1, 3, 2], natural language processing [4, 5, 6], generative models [8, 7, 9], visual question answering [10, 13], and abstract reasoning[14, 15, 16]. Deep learning, as a pivotal branch of machine learning, simulate the learning process of the human brain by establishing multilayered neural networks, thereby facilitating the learning and inference of intricate patterns within data[11]. Deep learning has found extensive applications to address diverse and intricate pattern recognition and reasoning problems[14, 15, 16].
Visual abstract reasoning problems occupy a pivotal position in the field of artificial intelligent. They involve identifying and comprehending hidden patterns, structures, and relationships within graphics, reflecting a deep cognitive process. Research on these problems has propelled advancements in artificial intelligent technology, fostering continuous optimization of key techniques such as feature extraction and modeling. The outcomes of this research have broad applications, spanning security, healthcare, and autonomous driving, among others. Furthermore, the capability of effective reasoning is a significant indicator of the intelligence level of AI systems. Addressing these problems poses challenges but also presents tremendous opportunities for breakthroughs, especially with the advent of deep learning techniques[12]. In summary, visual abstract reasoning problems are not only crucial to basic research but also hold the potential to drive technological progress and expand the application scope of artificial intelligent.
Visual abstract reasoning problems typically come in two forms: clustering reasoning and progressive pattern reasoning. For instance, Ravens Progressive Matrices (RPM)[14] serves as a benchmark for progressive pattern reasoning problems, while Bongard problems[15, 16] stand as a paragon for clustering inference, both encompassing learning requirements that span from perception to reasoning.
I-A Bongard-logo database
Bongard problems [15], contrasting with RPM (Raven’s Progressive Matrices) problems, are noteworthy exemplars of small-sample learning challenges. These problems typically present a series of images divided into two distinct groups: primary and auxiliary. The primary group comprises images that adhere to a specific set of rules defining an abstract concept, whereas the auxiliary group includes images that deviate from these rules in varying degrees. The intricacy of Bongard problems lies in the requirement for deep learning algorithms to precisely categorize ungrouped images into their corresponding groups based on subtle pattern recognition and abstract reasoning.
Within the domain of abstract reasoning, Bongard-logo[16] problems serve as a particular instantiation of Bongard problems, posing a significantly high level of reasoning difficulty. Each Bongard-logo problem consists of 14 images, with 6 images belonging to the primary group, 6 to the auxiliary group, and the remaining 2 serving as categorization options. These images are comprised of various geometric shapes, and the arrangements of these shapes serve as the basis for grouping. Figure 1 illustrates an exemplary Bongard-logo problem, showcasing the complexity and abstract nature of these challenges.
In figure 1, each Bongard problem is comprised of two distinct image sets: primary group A and auxiliary group B. Group A contains 6 images, with the geometric entities within each image adhering to a specific set of rules. Conversely, group B includes 6 images that deviate from the rules established in group A. The task at hand requires determining whether the images in the test set align with the rules dictated by group A. The level of difficulty associated with these problems varies depending on their structural complexity.
Figure 1: Bongard case
Furthermore, Bongard-logo problems are classified into three distinct conceptual categories: 1) Free Form problems (FF), in which each shape is composed of randomly sampled action strokes, potentially resulting in one or two shapes per image; 2) Basic Shape problems (BA), corresponding to the identification of a single shape category or combinations of two shape categories represented within given shape patterns; and 3) Abstract Concepts (HD), designed to assess a model’s proficiency in discovering and reasoning about abstract concepts such as convexity, concavity, and symmetry. These categorizations underscore the breadth and depth of reasoning abilities required to tackle Bongard-logo problems effectively.
I-B The RAVEN database
The RAVEN database[17], a subset of the broader RPM (Raven’s Progressive Matrices) problem set, typically comprises of 16 geometric entity-based images. These are structured into a problem statement consisting of 8 images and an answer pool containing the remaining 8 images. The objective for subjects engaging with this database is to select the appropriate images from the answer pool, enabling the completion of a 3x3 matrix within the problem statement. This matrix construction challenges individuals to discern specific abstract concepts embodied by the progressive patterns of geometric images arranged in a row-wise manner.
As exemplified in figure2, the construction of RAVEN problems adheres to a structured framework. Within this framework, human-defined concepts[11] inherent to geometric images, such as “shape” and “color”, undergo a process of artificial abstraction. This abstraction transforms these concepts into bounded, countable, and precise “visual attribute” values. Subsequently, the notion of “rule” is invoked to delineate the systematic progression of a circumscribed set of “visual attribute” values. It is noteworthy that certain visual attributes remain unconstrained by the “rule”, potentially introducing elements of interference in the reasoning processes of deep models.
Figure 2: RAVEN case
The RAVEN database showcases diversity with various sub-databases, including single-rule groups like center single, distribute four (G2x2), distribute nine (G3x3), and double-rule groups such as O-IC, U-D, L-R, and O-IG. Single-rule variations involve progressive attribute changes bound by one rule set, while double-rule variations add complexity with two independent rule sets.
I-C Pgm database
Pgm[18] and RAVEN problems exhibit a striking parallelism in their design philosophy, with both frameworks employing a problem statement consisting of 8 images and an answer pool supplemented by an additional 8 images. A unique facet of PGM problems is the expansion of the “rule” concept, which not only encompasses the progressive patterning of “visual attributes” horizontally within the matrix but also vertically. An illustrative example of a PGM problem is presented in the accompanying figure 3, highlighting its intricate nature.
Figure 3: PGM case
II Related work
This section is devoted to comprehensively presenting the significant research achievements related to the RPM (Raven’s Progressive Matrices) problem and the Bongard-Logo problem, as derived from past studies. These notable research milestones have played a pivotal role in driving forward the advancements in the field of artificial intelligence, specifically in relation to abstract reasoning capabilities. By thoroughly examining these milestones, we aim to foster a deeper understanding of the complexities and nuances associated with abstract reasoning challenges. Moreover, we hope to provide invaluable insights and guidance that can serve as a springboard for future research efforts, ultimately contributing to the continued growth and development of AI capabilities in tackling abstract reasoning tasks.
II-A RPM solver
Discriminative models designed for image reasoning tasks typically output a multidimensional vector, in which each dimension corresponds to the likelihood of selecting a particular graph from a set of candidate answers as the correct response. These models vary in their approach to learning and inference.
The CoPINet model[21] uses a contrast module to learn graph differences and an inference module to extract rules. LEN+teacher[22] employs a student-teacher approach for guided learning. DCNet[23] compares rule rows and columns with a dual contrast module. NCD[24] adopts unsupervised learning with pseudo-targets.
For subgraph analysis, SCL[25] sets up monitors to focus on specific visual attributes or rules. SAVIR-T[26] considers both within and between subgraph information. Recent studies show the benefits of decoupled visual features and symbolic attention for reasoning, as seen in PrAE[27].
ALANS[28] demonstrates strong generalization without prior rule knowledge. NVSA[29] uses holographic representations and ground-truth attributes for neural-symbolic reasoning.
RS-CNN[30] and RS-TRAN[30] combine various methods for RPM problems. RS-TRAN uses a Transformer-Encoder for multi-perspective inference, while RS-CNN integrates local and global information[41]. This paper introduces a new layer for RS-Tran, enhancing its efficiency and performance.
In summary, these models advance image reasoning research and pave the way for future progress.
II-B Bongard solver
In recent years, three main approaches have been employed to address Bongard problems: methods based on language feature models, methods based on convolutional neural network models, and generated datasets.
Methods based on language feature models[15]: somes researchers proposed a solution leveraging image-based formal languages, transforming image information into symbolic visual vocabulary, and subsequently tackling BP problems via symbolic languages and Bayesian inference. However, the limitation of this approach lies in its inability to directly apply to complex, abstract conceptual BP problems due to the need for reconstructing symbolic systems.
Methods based on convolutional neural network models[31]: some researchers constructed an image dataset consisting of simple shapes and utilized these images for pre-training to cultivate a feature extractor. Subsequently, image features from Bongard problems were extracted for image classification, determining whether test images conformed to the rules. Yun adopted a similar approach, initially pre-training with images containing visual features from BP problems to extract BP image features and then integrating an additional classifier for discrimination.
Methods based on Generated datasets[16]: some researchers employed fundamental CNNs, relation networks like WReN-Bongard, and meta-learning techniques, but their performance on the Bongard -Logo database was less than optimal.
These represent the main solutions targeted at Bongard problems in recent years, each with its own strengths and limitations.
II-C Capsule Network
Capsule Networks (CapsNets)[32] are a new type of deep neural network inspired by the “capsule” structures in the human brain, which process visual information efficiently. In CapsNets, neurons are grouped into capsules, and these networks use a dynamic routing algorithm to update weights, improving image feature recognition. Compared to traditional CNNs, CapsNets offer stronger feature representation and better performance. Each capsule detects a specific feature, outputting a vector where length indicates feature existence probability and direction encodes feature pose. CapsNets replace CNNs’ Max-Pooling with an agreement-based dynamic routing, adjusting network structure and weights based on input data, boosting performance and generalization. They also learn pose matrices for image representations, using a Gaussian Routing algorithm to weight high-dimensional representations, indicating pose consistency. CapsNets are mainly used in image recognition and have shown strong performance in tests.
II-D Transformer and Vision Transformer (ViT)
The Transformer [4] model, originally designed for machine translation tasks, is a encoder-decoder architecture that relies heavily on attention mechanisms, particularly self-attention. This mechanism enables the model to simultaneously consider other words in a sentence while processing a specific word, effectively capturing contextual relationships between them. On the other hand, Vision Transformer (ViT) [33] represents the application of the Transformer model in the image domain. It processes image data by dividing it into small patches, linearly embedding each patch into a vector, and then leveraging attention mechanisms to capture global contextual information for image recognition and processing. Both Transformer and Vision Transformer utilize self-attention mechanisms and encoder-decoder architectures, but they differ in the types of data they handle and the manner in which they process it.
II-E Sinkhron distance
The Sinkhorn distance[34] is a notion derived from the field of optimal transport theory, offering a computationally efficient alternative to the classical Wasserstein distance[35]. At its core, it approximates the cost of transporting mass from one probability distribution to another by relaxing the constraint of mass preservation through a regularization term. This regularization, typically parameterized by a scalar value known as the entropy regularization parameter, allows for a smoother, more tractable optimization problem.
II-F Discrete-Time Inference and Establishment of Contextual Probability Models
The theory of isochronous constraint reasoning [36] has brought profound insights to the design of reasoning problem solvers. By integrating logical reasoning from this theory, along with optimization strategies and discrete-time system analysis, we can more precisely understand and predict complex logical relationships in reasoning problem. Additionally, drawing on the experience of isochronous constraint reasoning in addressing nonlinear problems, and adopting data-driven techniques, we can build an efficient and accurate reasoning problem solver[37], effectively tackling challenges in visual entities and elements sequence reasoning.
The contextual analysis[38] method provides significant insights into the design of abstract reasoning problem solvers. By deeply analyzing the contextual relationships between graphics, the solver can more accurately grasp the logical connections and patterns of change among them. Specifically, this method urges consideration of the position and role of graphics in the overall sequence, leading to a more comprehensive understanding of each graphic’s significance. Encouraging a focus on the mutual influence among graphics rather than viewing each in isolation, the solver’s design benefits from capturing and utilizing contextual information, thereby enhancing problem-solving accuracy and efficiency, especially when dealing with complex and nonlinear graphic relationships. In conclusion, the contextual analysis approach serves as a powerful thinking tool for the design of visual abstract reasoning problem solvers, facilitating the development of smarter and more efficient algorithms.
Attention-based visual models[39] demonstrate significant superiority by focusing on critical information within images, thereby enhancing processing efficiency and accuracy. These models enable a sharper focus on pertinent data while adapting to multi-modal data relationships, resulting in their widespread application across various domains. It also serves as the underpinning of our methodological approach.
In light of the challenges and opportunities associated with modeling system problem spaces using probabilistic models[40], this paper opts to leverage deep learning algorithms to construct a deep-learning-based probabilistic model aimed at addressing visual abstract reasoning problems.
III Methodology for Bongard-Logo
The Bongard-logo[15] and Raven’s Progressive Matrices (RPM)[14] tasks are both paradigms of abstract reasoning that require the extractor’s ability in unveiling precise concepts hidden within abstract representations. These concepts are indicative of more elevated abstractions, surpassing rudimentary ideas such as “pixel configuration patterns”. Frequently, these elevated abstractions encapsulate diverse degrees of human-centered pre-existing knowledge knowledge, including elements like form, dimension, hue, spatial orientations, concave and convex contours of shapes, and shape completion. Using distinct querying approaches, the Bongard-logo and RPM tasks endeavor to evaluate the extractor’s expertise in unveiling and assimilating these abstract concepts.
The methodology presented in this section targets clustering reasoning problems within visual abstract reasoning, attempting to solve them by modeling a probabilistic model. The Bongard-Logo problem stands as a renowned example of clustering reasoning, drawing significant attention due to its exceptional reasoning difficulty and the current lack of progress in solving it. This paper aims to achieve promising results on the Bongard-Logo problem, demonstrating the potential of probabilistic models in solving clustering reasoning tasks.
III-A A method for solving the bongard-logo problem based on sinkhorn distance (SBSD)
In the Bongard-Logo problem, we denote $x_{ij}$ as the $j$-th image in the $i$-th Bongard-Logo problem, where $i\in[1,n]$ signifies the question number, with $n$ being the total number of questions. Specifically, $\{x_{ij}\}_{j=1}^{6}$ represents images in the $i$-th primary group, while $\{x_{ij}\}_{j=8}^{13}$ represents images in the $i$-th auxiliary group. Additionally, $x_{i7}$ represents the test image to be potentially assigned to the $i$-th primary group, and $x_{i14}$ represents the test image to be potentially assigned to the $i$-th auxiliary group. we let the distribution of the primary group (positive examples) within a Bongard-Logo question be denoted as $p_{i}(x|y)$, and the distribution of the auxiliary group (negative examples) be denoted as $q_{i}(x|y)$. Here, y represents the reasoning type condition of the question, where $y\in\{ff,ba,hd\}$. We are committed to developing a deep learning algorithm that can induce distributions $p_{i}(x|y)$ and $q_{i}(x|y)$ with low cross-measure between them.
Directly deriving the image distributions $p_{i}(x|y)$ and $q_{i}(x|y)$ of Bongard-Logo without dimensionality reduction and computing the cross measure between these two distributions is quite challenging. Moreover, such an approach would also result in a waste of resources and time. Therefore, we aim to devise a deep learning algorithm, denoted as $f_{\theta}(z|x)$, which can map samples $x_{ij}$ to latent variables $z_{ij}$. We hope that $f_{\theta}(z|x)$ can serve as a distribution transformation function, and by measuring the distribution of the latent representation, we can address the Bongard-Logo problem. Accordingly, the key objective is to ensure that the encoder $f_{\theta}(z|x)$ possesses a specific image encoding capability, which necessitates its ability to produce encoded latent variables that demonstrate specific characteristics. Specifically, the distribution of latent variables within the primary group (represented as $p_{i}^{\prime}(z|y)$) and the distribution of latent variables within the auxiliary group (represented as $q_{i}^{\prime}(z|y)$) should exhibit very low cross-measure. The Kullback-Leibler (KL) divergence, a measure used to quantify the similarity between distributions, can be employed as an optimization objective for $f_{\theta}(z|x)$[7]. The formula for calculating KL divergence is provided below.
$KL(P||Q)=\sum P(x)\log\left(\frac{P(x)}{Q(x)}\right)$ (1)
However, due to the characteristics of small sample learning problems, it is challenging to estimate the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ of Bongard-Logo image representations. Estimating the specific forms of distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ poses an even greater challenge. Additionally, the limitations of the Kullback-Leibler (KL) divergence make it difficult to directly optimize for these conditions and obtain a highly effective deep model.
Hence, we endeavored to utilize baseline models, specifically ResNet18, for the purpose of encoding images within the Bongard-Logo paradigm, denoted as $\{x_{ij}\}_{j=1}^{14}$, into meaningful representations, designated as $\{z_{ij}\}_{j=1}^{14}$. These representations, in the context of a Bongard-Logo question, can be interpreted as samples drawn from the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$. Subsequently, constraints were imposed on the Sinkhorn distance between $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ through these representations. This methodology was designed primarily because the Sinkhorn algorithm computes the distance without requiring explicit knowledge of the distributions’ specific forms, making it highly suitable for scenarios where the forms of $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ remain unknown.
Specifically, the representations within the primary group, denoted as $\{z_{ij}\}_{j=1}^{7}$, were randomly partitioned into two distinct subgroups: $\{z_{it}\}$ and $\{{z}_{i\tilde{t}}\}$. By utilizing the Sinkhorn distance between these subgroups, constraints were imposed to ensure that the representations within the primary group $\{z_{ij}\}_{j=1}^{7}$ adhered to a consistent, albeit unknown, distribution. Analogously, similar constraints were introduced between the distributions represented by $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$, leveraging representations from both the primary and auxiliary groups to minimize the cross-measure.
The aforementioned constraints can be explicitly articulated through a mathematical loss function.
(${\ell}(\{z_{ij}\}_{j=1}^{14})$
$=-\log\frac{{{e^{-D(\{z_{it}\},\,\{{z}_{i\tilde{t}}\})}}}}{{{e^{-D(\{z_{it}\},\,\{{z}_{i\tilde{t}}\})}}+{e^{-D(\{z_{ij}\}_{j=1}^{7},\,\{{z}_{ij}\}_{j=8}^{14})}}}}$ (2)
Let $D(a,b)$ denote the calculation of the Sinkhorn distance between the distributions associated with sets $a$ and $b$. Additionally $\{z_{it}\}\cup\{{z}_{i\tilde{t}}\}=\{{z}_{ij}\}_{j=1}^{7}$ and $\{z_{it}\}\cap\{{z}_{i\tilde{t}}\}=\emptyset$. Despite our endeavors, this approach did not yield satisfactory results. We hypothesize that this was primarily due to the limited sample size within a single Bongard-logo problem, which, when collectively treated as sampling outcomes, may not meet the prerequisite sampling requirements of the Sinkhorn distance. An illustrative diagram depicting this process can be found in Figure 4.
Figure 4: Feedforward process of SBSD
In summary, SBSD represents our initial and most direct attempt to model the Bongard-Logo problem as a probabilistic framework. SBSD treats the representations of images within the primary group as samples from the primary group distribution $p_{i}^{\prime}(z|y)$, and the auxiliary group follows the same approach. By utilizing these samples, the Sinkhorn distance is employed to calculate and represent the distance between the primary $p_{i}^{\prime}(z|y)$ and auxiliary distributions $q_{i}^{\prime}(z|y)$. This distance can be used to solve the Bongard-Logo problem.
III-B PMoC(Probability Model of Concept)
After conducting experiments, we discovered that SBSD had achieved limited progress in reasoning accuracy. However, this paper aims to make more contributions and breakthroughs. Given the difficulties associated with utilizing the SBSD approach, which involves optimizing deep networks based on the distances between distributions, this paper proposes the design of the PMoC (Probability Model of Concept) network. PMoC explores an alternative approach by shifting the constraint conditions. The training objective shifts from focusing on the deep model’s ability to disentangle distribution $p_{i}^{\prime}(z|y)$ and distribution $q_{i}^{\prime}(z|y)$ to emphasizing its capacity to compute the probability of a sample belonging to the distribution $p_{i}^{\prime}(z|y)$. For this purpose, we assume that $p_{i}^{\prime}(z|y)$ follows a multivariate Gaussian distribution.
From a macroscopic perspective, our PMoC framework consists of two components: module $f(z|x)$ and module $g(\mu,\sigma^{2}|z)$. We leverage the $f(z|x)$ module to map Bongard-Logo images into latent representations, and the $g(\mu,\sigma^{2}|z)$ module to deduce a Gaussian distribution based on the representations of samples from the primary group. Subsequently, we compute the probabilities of representations from the auxiliary group and the test samples belonging to this distribution. Finally, we employ a cross-entropy loss function to constrain these probabilities, optimizing both the mapping capability of the $f(z|x)$ module and the distribution computation ability of the $g(\mu,\sigma^{2}|z)$ module.
Specifically, we utilize a deep network $f_{\theta}(z_{ij}|x_{ij})$ (serving as the $f(z|x)$ module) to transform the Bongard-Logo images $\{x_{ij}\}^{14}_{j=1}$ into their corresponding latent representations $\{z_{ij}\}^{14}_{j=1}$. Then, another deep network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ (acting as the $g(\mu,\sigma^{2}|z)$ module) is employed to derive an optimizable multivariate Gaussian distribution $p^{\prime}_{i}(z|\{z_{ij}\}^{6}_{j=1},\omega)$ based on the latent representations of the primary group samples $\{z_{ij}\}^{6}_{j=1}$. Next, we compute the probabilities of each latent representation $z_{ij^{\prime}}$ (where $j^{\prime}$ ranges from 7 to 14) belonging to this distribution, denoted as the set $\{p^{\prime}_{i}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$. In essence, the network $f_{\theta}(z_{ij}|x_{ij})$ functions as an image encoder, mapping input images to their latent representations, while $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ serves as a logical distribution fitter, calculating the distribution of latent representations and modeling it as a Gaussian distribution. To this end, we adopt a Convolutional Neural Network (CNN) to fit $f_{\theta}(z_{ij}|x_{ij})$ and a Transformer-Encoder to fit $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$. Further details regarding the design are provided below. This framework is visualized in Figure 5.
Figure 5: Framework of PMoC.
Regarding the network implementation of $f_{\theta}(z_{ij}|x_{ij})$, we present an approach that combines ResNet18 with Transformer-Encoder, employing this integration as the architecture for the function $f_{\theta}(z_{ij}|x_{ij})$. In detail, we encode the Bongard-Logo problem images $x_{ij}$ into feature maps $h_{ij}\in R^{h\times w\times d}$ using ResNet18. Subsequently, we calculatethe self-attention results $z^{\prime}_{ij}\in R^{n\times d}$ among all receptive fields within the feature maps $h_{ij}$. This computation enables the $n$ attention outputs $z^{\prime}_{ij}$ to capture global information of the image in varying degrees, ensuring that each output retains some level of contextual understanding. We encoded Bongard-Logo images from multiple scales and perspectives using the aforementioned method. We denote each individual perspective within $z^{\prime}_{ij}$ as $z_{ij}\in R^{d}$. Since $g_{\omega}(\cdot|\cdot)$ in this paper treats each perspective within $z^{\prime}_{ij}$ equally and in parallel, we do not assign different notations to individual perspectives within $z^{\prime}_{ij}$ and uniformly denote them as $z_{ij}$. The feedforward process of the network $f_{\theta}(z_{ij}|x_{ij})$ is illustrated in the figure 6.
Figure 6: Feedforward process of $f_{\theta}(z_{ij}|x_{ij})$
The network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ is much more of its name. It encodes the mean and variance of the $i$-th primary group distribution based on the latent representations of the $i$-th primary group samples, and $i$ still represents the problem index in Bongard-Logo. Specifically, we uniformly process each perspective $z_{ij}$ within $z^{\prime}_{ij}$. Under a single perspective, we treat $\{z_{ij}\}^{6}_{j=1}$ as six tokens and, after appending two optimizable vectors along with positional embedding, input them into $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ with a Transformer-Encoder backbone for computing attention results. We extract two optimizable vectors from the attention results, serving as the mean $\mu_{i}\in R^{d}$ and logarithm of variance $\log(\sigma_{i}^{2})\in R^{d}$ for a multivariate Gaussian distribution. The encoding distribution pattern of module $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ follows a similar approach to that of VAE[7]; however, it does not incorporate variational inference in its design. The aforementioned process parameterizes $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and outlines the computation of $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ and this process can be expressed as Figure 7.
Figure 7: Parameterizes process of $p_{i}^{\prime}(z|y)$
We feed 8 vectors $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ into this distribution for calculating 8 logical probabilities $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$ in one perspective. In this paper, we calculate the final probability ${p}_{i}^{\prime}(z^{\prime}_{ij^{\prime}})$ by averaging the logical probabilities ${p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)$ obtained from all perspectives. The calculation process of the final probability can be expressed as Figure 8.
Figure 8: The calculation process of the final probability
By utilizing the cross entropy loss function, we constrain the final probability ${p}_{i}^{\prime}(z^{\prime}_{i7})$ to approach 1 while ensuring other probabilities $\{{p}_{i}^{\prime}(z^{\prime}_{ij})\}_{j=8}^{14}$ are close to 0. Figure 9 illustrates in detail the forward process of the PMoC. However, this approach did not yield satisfactory results.
Figure 9: Feedforward process of PMoC.
We speculate that the reason for the ineffectiveness of the aforementioned method lies in the unknown true form of the distribution $p_{i}^{\prime}(z|y)$, which may be a complex mixed Gaussian distribution. Therefore, it may not be reasonable to parameterize $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and compute the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Additionally, the formula for calculating the logical probability of a vector $z$ belonging to the parameterized distribution is provided as follows:
$\log(p(z,\mu,\sigma))=\frac{{(z_{m}-\mu_{m})^{2}}}{{2\sigma^{2}_{m}}}-\log(|\sigma_{m}|)-\frac{1}{2}\log(2\pi)$ (3)
Where $m$ represents the dimension index of the vector. $\mu$ represents the mean of the multivariate Gaussian distribution, $\sigma$ represents the standard deviation of the multivariate Gaussian distribution, and $z$ is the representation of some unknown data. It is evident that optimizing the network using the given formula (3) as the loss function presents challenges, as optimizing the above equation can be indirectly viewed as optimizing the Mean Squared Error (MSE) loss between vectors. Moreover, these vectors have not undergone normalization or squashing, making it difficult to optimize the network from the perspective of implementation.
To counter such ineffectiveness, we decide to train a deep network to directly fit probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. We argue that, rather than laboriously analyzing and fitting a distribution, it is more efficient to directly compute the probability of a sample belonging to that distribution. Inevitably the network’s formulation has been transformed from $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ to $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$, thereby shifting the problem. Different from previous approaches $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ that fit the distribution $p_{i}^{\prime}(z|y)$, this paper employs a Transformer-Encoder to fit the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$ directly, which enables the direct computation of the probability that $z_{ij^{\prime}}$ belongs to the distribution $p_{i}^{\prime}(z|y)$.
For the implementation on the new form of $g_{\omega}(\cdot|\cdot)$, the function $g_{\omega}(\cdot|\cdot)$ maintains the egalitarian approach in processing each perspective $z_{ij}\in R^{d}$ within the outcomes $z^{\prime}_{ij}\in R^{n\times d}$ encoded by $f_{\theta}(z_{ij}|x_{ij})$. We amalgamate $\{z_{ij}\}^{6}_{j=1}$ with one of the pending representations $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, resulting in a set of seven tokens. Following the embedding of positional encoding, these tokens are then systematically fed into a Transformer-Encoder, aimed at yielding 7 attention results. All 7 attention results are mapped to 7 probabilities by a MLP. The mean of the mapped outcomes is considered as the logical probability ${p}_{i}^{\prime}(z_{ij^{\prime}}|y)$ corresponding to the specified perspective. This constitutes the feedforward process of the $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$. This feedforward process can be expressed as Figure 10.
Figure 10: Feedforward process of $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$.
Each pending representation $z_{ij^{\prime}}$ within $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ receives a corresponding probability output $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Each logical probability stored in $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$ is obtained based on a single perspective. By averaging the logical probabilities across all perspectives, we derive the final probability values $\{p_{i}^{\prime}(z^{\prime}_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$. The new calculation process of the final probability can be expressed as Figure 11.
Figure 11: The new calculation process of the final probability
In this paper, we still optimize the model using the cross-entropy loss function. It is worth mentioning that the essence of the final version of PMoC lies in its function as a distribution distance metric, thereby necessitating the satisfaction of Lipschitz continuity. To ensure this, we have imposed spectral normalization constraints on the parameters of PMoC. Figure 12 visually depicts the process of PMoC.
Figure 12: Feedforward process of PMoC 2.0 ver.
In summary, this paper makes two compromises for PMoC. Firstly, we acknowledge that compared to $p_{i}^{\prime}(z)$, $q_{i}^{\prime}(z)$ is more likely to be a complex distribution. For instance, $p_{i}^{\prime}(z)$ may follow a Gaussian distribution, while $q_{i}^{\prime}(z)$ may adhere to a mixed Gaussian distribution. Therefore, estimating these two distributions and utilizing them to solve the Bongard-Logo problem is challenging to achieve excellent performance. Accordingly, we shift our focus to estimating the distribution $p_{i}^{\prime}(z)$ and utilize the probability of a representation belonging to this distribution to determine whether it is a primary or auxiliary group representation. Secondly, due to the challenges associated with optimizing the model from the perspective of estimating $p_{i}^{\prime}(z)$, this paper opts to directly compute the probability of a representation belonging to the distribution $p_{i}^{\prime}(z)$.
Despite the apparent disadvantage of probabilistic clustering reasoning solvers, we persist in designing and developing a Bongard-Logo solver in the form of a probabilistic model. Our aim is not to rival the most advanced solutions available, but rather to contribute to the diversity of deep learning research outcomes. We are motivated by the goal of fostering innovation and exploration beyond the mainstream, encouraging a wider range of approaches and perspectives in the field. We are confident that a probabilistic clustering approach can offer more intuitive probability assessments for clustering decisions, which could be critically important in certain analogous fields.
IV Methodology for abstract reasoning
Previous research has demonstrated limited success in learning abstract concepts related to positional relationships among entities, particularly when addressing the Bongard-Logo and RPM problems. This limitation is evident in challenges posed by the OIG and D-9 problem sets from RAVEN, as well as the ff and hd issues of Bongard-Logo and the Neurtal problem in PGM.
In this section, we introduce the Pose-Transformer, an enhancement to the Transformer-Encoder framework that incorporates the pose matrix from the Capsule Network[32]. This integration implicitly enhances the Transformer-Encoder’s ability to discern positional relationships among local representations, thereby bolstering its learning capabilities on visual abstract reasoning problem. Furthermore, we have lightweighted Pose-Transformer into Straw-Pose-Transformer, which achieves a lightweight version of the former while maintaining its effectiveness. We present and delve into the intricacies of our proposed Pose-Transformer, discussing its potential to revolutionize learning in tasks that involve complex positional relationships through experiments.
IV-A Pose-tansformer
We delve into the intricacies of the Transformer-Encoder framework, particularly when it processes logical inputs to generate attention-weighted outputs. During this process, all output tokens become associated with varying degrees of global information derived from the logical inputs[26, 33]. Operating on a multi-head self-attention mechanism, the Transformer-Encoder segments the input tokens into multiple sub-vectors and computes their self-attention results in parallel[33]. From an informational standpoint, we posit that these segmented sub-vectors possess lesser information compared to the holistic tokens encoding global information, effectively rendering them as representations of localized information. Consequently, the relative positional relationships among these localized informational units warrant a summarization and analysis.
Notably, in Capsule Networks[32], a set of pose mapping matrix is set to map local information features into pose vectors, and the Gaussian routing algorithm is employed to calculate the pose consistency between these pose vectors, thereby gauging the relative positional relationship among local features[32]. This paper proposes that a similar pose mapping process could be integrated into the Transformer-Encoder, enabling it to be more sensitive to the relative positional relationships among local features when handling visual tasks. Therefore, we have designed Pose-Transformer and have high expectations for its performance on visual abstract reasoning problems.
Specifically, we have designed a novel fundamental module termed “pose matrix embedding”. When logical inputs $\{\tilde{V}_{i}|i\in[1,n],\tilde{V}_{i}\in R^{d}\}$ pass through a single Transformer-Encoder layer, this paper segments the $n$ output tokens $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$, attention results, into $m$ local vectors $\{H_{ij}|i\in[1,n],j\in[1,m],H_{ij}\in R^{d/m}\}$ of the attention head size. A set of mapping matrices $\{W_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],W_{ijk}\in R^{(d/m)\times d}\}$ is established for each local vector, with the number of matrices $k$ within a set being identical to the token count $n$ of the logical input. These mapping matrices facilitate the transformation of local vectors into pose vectors $\left(H_{ij}\times W_{ijk}\right)$. We compute the entire set of pose vectors $\{P_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],P_{ijk}\in R^{d}\}$ resulting from the mapping of each local vector. By summing and squashing these pose vectors and embedding them to $\{V_{i}|i\in[1,n],Vi\in R^{d}\}$, we obtain the logical output $\{O_{k}|k\in[1,n],O_{k}\in R^{d}\}$ of the foundational block of Pose-Transformer. The squashing function is formulated as shown in the equation below:
$\text{squash}(\mathbf{v})=\frac{||{\mathbf{v}}||^{2}}{1+||{\mathbf{v}}||^{2}}\cdot\frac{\mathbf{v}}{||{\mathbf{v}}||}$ (4)
where the operator |||||$\cdot$||| | ⋅ | | is $L_{2}$ norm. It is worth mentioning that the squash function, $squash(\cdot)$, is employed in Capsule Networks to squash pose vectors for computing pose consistency among all local representations. Similarly, in the context of Pose-Transformer, the squash function continues to serve its purpose and retain its significance, contributing to the overall architecture’s efficacy. The feedforward process of a single pose matrix embedding fundamental block is illustrated in the figure 13. The entire calculation process of “pose matrix embedding” can be expressed as fallows:
[…]
$\{H_{ij}|\,j\in[1,m]\}$ $=$ $\text{segment}(V_{i})$ (6)
$P_{ijk}$ $=$ $H_{ij}\times W_{ijk}$ (7)
$O_{k}$ $=$ $\,(V_{i}|i=k)+\text{squash}(\sum_{i=1}^{n}\sum_{j=1}^{m}{P_{ijk}})$ (8)
In the formula, the operator “segment” represents the aforementioned segment process. The process of feeding the input vector set into the Transformer-Encoder to obtain the attention results is denoted as “T-Encoder”.
Figure 13: The feedforward process of a single pose matrix embedding block
This fundamental block is followed by a feedforward block, forming the basic building block of the Pose-Transformer. The design of the feedforward block is consistent with that in the Transformer-Encoder architecture. Notably, the proposed fundamental block is dimension-preserving, allowing for the stacking of multiple blocks to adapt to different scales of visual reasoning problems. The Pose-Transformer with $N$ logical layers is depicted in the figure 14. When $N=1$, the Pose-Transformer collapses into a regular Transformer-Encoder . In Capsule Networks, similarity weights are distributed across higher-level capsules and computed in a bottom-up fashion, whereas attention weights in the Transformer-Encoder are distributed over lower-level representations and computed top-down. In the Pose-Transformer with $N>1$, similarity weights for local representations can be calculated at multiple levels and scales from the outside in. This is ensured by the fact that each logical layer in the Pose-Transformer is succeeded by a standard Transformer-Encoder layer, which can be present in the subsequent logical layer or the terminating layer, responsible for recalculating the similarity weights.
Figure 14: The architecture of the Pose-Transformer
The Pose-Transformer diverges from the dynamic routing mechanism of Capsule Networks, which iteratively computes representation similarity weights. Instead, it employs an attention mechanism to perform the task of the routing. For the complete computation of similarity weights, the Pose-Transformer terminates with a standard Transformer-Encoder layer, which serves as the final stage for calculating similarity weights on the pose vectors.
IV-B Straw Pose-Tansformer
Pose-Transformer, during deployment, exerts considerable pressure on computational resources. The establishment and computation of the mapping matrices, denoted as $\{W_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],W_{ijk}\in R^{(d/m)\times d}\}$, incur significant resource overhead. Motivated by this challenge, this paper explores the lightweighting of Pose-Transformer. The crux of Pose-Transformer lies in computing compressed pose vectors and subsequently employing them to modify and refine the logical inputs $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to some extent. The computation of the logical outputs in Pose-Transformer’s pose matrix mapping follows the methodology outlined below:
$O_{k}=V_{i}+squash\left(\sum_{i}^{n}\sum_{j}^{m}{H_{ij\_}}\times W_{ijk}\right)$ (9)
It is worth noting that the placeholders marked with “_” in the formula represent numbering for on-demand broadcasting. Multiple levels of summation and mapping can incur significant computational overhead. By designing parameter sharing, we aim to eliminate partial logical summation and mapping, thereby achieving the purpose of reducing computational cost. From a logical hierarchy perspective, only the logical mapping at the $i$-th level possesses the potential for indentation, whereas the other logical levels exhibit stringent mapping requirements. In other words, this paper posits that the pose matrices between equivalent local vectors across different tokens merit weight sharing. The planned target expression of pose matrix mapping after indentation is:
$O_{k}=V_{i}+squash\left(\sum_{i}^{n}\sum_{j}^{m}{H_{j\_}}\times W_{jk}\right)$ (10)
This paper aims to simplify the Pose-Transformer in a manner consistent with the aforementioned expression. We treat the mapping matrix $W_{ijk}$ as a product of the mapping matrix $\tilde{W}_{jk}$ multiplied by $i$ different coefficients, specifically expressed as $W_{ijk}\tilde{=}w_{i}\cdot\tilde{W}_{\_jk}$. This transformation significantly reduces the number of mapping matrices involved. We argue that it is more reasonable to combine $w_{i}$ with $H_{ij\_}$ and treat them as a weighted sum of $H_{ij\_}$, rather than treating $w_{i}$ as a set of optimizable preliminary coefficients for mapping matrices. Furthermore, treating this process as computing an attention-based weighted summation for $H_{ij\_}$ is even more elegant and effective. This solution aligns well with the formula presented:
$O_{k}=$ $V_{i}+squash\left(\sum_{i}^{n}\sum_{j}^{m}{H_{ij\_}}\times W_{ijk}\right)$ (11)
$\tilde{O}_{k}=$ $V_{i}+squash\left(\sum_{i}^{n}\sum_{j}^{m}{H_{ij\_}}\times(w_{i}\cdot\tilde{W}_{\_jk})\right)$
$=$ $V_{i}+squash\left(\sum_{j}^{m}(\sum_{i}^{n}w_{i}\cdot{H_{ij\_}})\times\tilde{W}_{\_jk}\right)$
$=$ $V_{i}+squash\left(\sum_{j}^{m}{\tilde{H}_{\_j\_}}\times\tilde{W}_{\_jk}\right)$
$=$ $V_{i}+squash\left(\sum_{j}^{m}{\tilde{H}_{j\_}}\times\tilde{W}_{jk}\right)$
$\tilde{=}$ $V_{i}+squash\left(\sum_{j}^{m}Attention({H_{ij\_}})\times\tilde{W}_{\_jk}\right)$ (12)
Notably, the aforementioned approach necessitates additionally computing attention-based weighted summation between the standard Transformer-Encoder and pose mapping. In other words, between equations (5) and (6) lies the necessity to incorporate an additional computational step, which prolongs the training time, contradicting the intended purpose. This paper utilizes a Transformer-Encoder with an additional mask mechanism to integrate the Transformer encoding process with the additional attention-based weighted summation process.
Specifically, this paper introduces an optimizable vector, denoted as $\tilde{V}_{0}\in R^{d}$, for the embedding of logical inputs $\{\tilde{V}_{i}|i\in[1,n]\}$. After the embedded logical inputs $\{\tilde{V}_{i}|i\in[0,n]\}$ propagate through a single Transformer-Encoder layer, while applying a mask to the self-attention weights at the position corresponding to $\tilde{V}_{0}$ to prevent its value from contributing to the attention results of the Transformer-Encoder, they produce an updated set of vectors $\{V_{i}|i\in[0,n]\}$ representing the attention-weighted outputs. This process effectively entails the concurrent computation of two procedures: the self-attention calculation within $\{\tilde{V}_{i}|i\in[1,n]\}$ and the cross-attention computation between $\tilde{V}_{0}$ and $\{\tilde{V}_{i}|i\in[1,n]\}$. Accordingly, $\{V_{i}|i\in[1,n]\}$ represents the self-attention outcome of $\{\tilde{V}_{i}|i\in[1,n]\}$, while $V_{0}$ corresponds to the weighted sum of the vectors within $\{\tilde{V}_{i}|i\in[1,n]\}$, where the weights are determined by their attention similarity scores with $\tilde{V}_{0}$. This masked attention can be expressed as Figure 15. In Figure 15, we show a simplified case with only three vectors in $\{\tilde{V}_{i}\}$ for clarity.
Figure 15: Masked attention
We replace the tedious sequential process with this masked attention mechanism, thereby streamlining the computational workflow.
Due to the inherent self-attention mechanism and masking, the vector $\tilde{V}_{0}$ assimilates information from the logical inputs $\{\tilde{V}_{i}|i\in[1,n]\}$, transforming into global vector $V_{0}$. Within this global vector, we establish a collection of mapping matrices $\{\tilde{W}_{jk}|j\in[1,m],k\in[1,n],\tilde{W}_{jk}\in R^{(d/m)\times d}\}$ for each attention head $\{H_{0j}|j\in[1,m],H_{0j}\in R^{(d/m)}\}$ in it.
The pose vectors $\{p_{jk}|j\in[1,m],k\in[1,n],p_{jk}\in R^{d}\}$ is achieved by multiplying the attention heads with the mapping matrices. These pose vectors, after summation and squashing, are added to the vector set $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to obtain the output of new pose matrix mapping block $\{O_{i}|i\in[1,n],O_{i}\in R^{d}\}$. This process outlines the computation process of the pose embedding matrix block within the “Straw-Pose-Transformer” architecture. The entire calculation process of new “pose matrix embedding” can be expressed as fallows:
[…]
$\{H_{0j}|\,j\in[1,m]\}$ $=$ $\text{segment}(V_{0})$ (14)
$P_{jk}$ $=$ $H_{0j}\times\tilde{W}_{jk}$ (15)
$O_{k}$ $=$ $\,(V_{i}|i=k)+\text{squash}(\sum_{j=1}^{m}{P_{jk}})$ (16)
In the formula, the process of the Masked Transformer-Encoder to obtain the output is denoted as “Masked T-Encoder”. In the Straw-Pose-Transformer architecture, the outputs from each pose matrix embedding block, denoted as $\{O_{k}|k\in[1,n]\}$, undergo a concatenation process with the global vector $V_{0}$. This modified set $\{V_{0},O_{k}|k\in[1,n]\}$, now represented as $\{\tilde{V}_{i}|i\in[0,n]\}$ , is then propagated into the subsequent logical layer for further processing. However, if the subsequent logical layer is the final layer, the global vector $V_{0}$ is no longer appended.
In this way, we can reconstruct the pose matrix embedding process according to the formula (10). The feedforward process of new pose matrix embedding is illustrated in the figure 16. And figure 17 illustrate the structure of the Straw-Pose-Transformer.
Figure 16: The feedforward process of a new pose matrix embedding block
Figure 17: The architecture of the Straw-Pose-Transformer
In summary, our lightweighting strategy for Pose-Transformer involves the insertion of additional cross-attention computations between the original Transformer-Encoder calculations and the pose matrix mapping computations. While this approach reduces the number of parameters, it also elongates the training time, which is an undesirable trade-off. Therefore, in this study, we have designed a masked Transformer-Encoder that obviates the need for explicitly inserting cross-attention computation processes, aiming to achieve a more efficient and streamlined model.
V combination of Pose-Transformer and baselines
In addressing the Bongard-Logo problem, this paper introduces a novel baseline, PMoC. To demonstrate the adaptability of the Straw-Pose-Transformer in solving the Bongard-Logo problem, the Transformer-Encoder component within PMoC is replaced with the Straw-Pose-Transformer. Furthermore, given the remarkable performance achieved by the predecessor RS-Tran[30] in solving the RPM problem, this paper does not design an additional baseline network to prove the superiority of the Pose-Transformer in addressing the RPM problem. This paper aims to demonstrate the advantage of Pose-Transformer over the standard Transformer-Encoder for RPM problems by achieving improved accuracy through the simple replacement of the Transformer-Encoder in RS-Tran with Pose-Transformer.
The RS-Tran[30] primarily consists of Transformer-Encoder structures, and in this study, we replace it with the Pose-Transformer and Straw-Pose-Transformer and conduct experiments. The modified RS-Tran is renamed as RS-Pose-Tran and RS-Straw-Pose-Tran. Due to the fundamental modifications made to the RS-Tran architecture, RS-Pose-Tran and RS-Straw-Pose-Tran can be regarded as models designed based on our novel structure, incorporating inductive biases inherent in the RPM problem[41].
VI Experiment
In this section, we conduct experiments on the model proposed in this paper.
VI-A Experiment on Bongard-Logo
All our experiments on Bongard-Logo are implemented in Python using the PyTorch[42] framework. In the experiments conducted for PMoC and SBSD, the Adam[43] optimizer was selected with a learning rate set to 0.001, a learning rate decay of 0.995, a weight decay of 0.0001, and a batch size of 40 for PMoC and SBSD. Specifically, in the experiment combining PMoC with Straw-Pose-Transformer, the batch size was set to 20.
We conducted experiments on the Bongard-logo dataset, utilizing SBSD, PMoC, and the integration of PMoC with Straw-Pose-Transformer. The experimental results are documented in Table I. Due to the propensity of the serial convolutional scanning mechanism and attention mechanism to experience attention collapse when combined, we adopted the parameters of the SBDC model as pre-trained parameters for the PMoC convolutional layer. Notably, under the experimental conditions of this study, which involved the use of four A100 Graphics Processing Units (GPUs), the integration of PMoC with Pose-Transformer posed significant deployment challenges. This served as one of the key reasons and motivations behind our decision to develop a lightweight approach for the Pose-Transformer in this paper.
TABLE I: Reasoning Accuracies of PMoC on Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
SNAIL 59.2 56.3 60.2 60.1 61.3
ProtoNet 73.3 64.6 72.4 62.4 65.4
MetaOptNet 75.9 60.3 71.6 65.9 67.5
ANIL 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
Meta-Baseline-MoCo 81.2 65.9 72.2 63.9 64.7
WReN-Bongard 78.7 50.1 50.9 53.8 54.3
SBSD 83.7 75.2 91.5 71.0 74.1
PMoC 92.0 90.6 97.7 77.3 76.0
PMoC+Straw-Pose-Transformer 93.2 91.4 98.0 77.5 76.3
PMoC+Pose-Transformer Deployment is not feasible on our device
In Table I, it is evident that addressing the Bongard-Logo problem using a probabilistic model approach, namely PMoC, is both straightforward and effective. Furthermore, SBSD exhibits stronger outcomes when compared to previous models. Additionally, the integration of Straw-Pose-Transformer has contributed to enhancing the inference accuracy of PMoC.
Finally, we postulate that the significance of PMoC, a network primarily structured around the Transformer-Encoder, may be further emphasized when dealing with larger datasets. ViT [33] explicitly explores the importance of data volume for Transformer-based visual models, whereas RS-Tran [30] focuses on the significance of data quantity in addressing visual abstract reasoning challenges. Therefore, this paper attempts to conduct experiments on an expanded Bongard-Logo dataset to explore the further capabilities of PMoC and (Straw) Pose-Transformer. However, due to the limited availability of additional Bongard-Logo data for expanding the dataset, this paper uses the data augmentation techniques to achieve the objective. To expand the Bongard-Logo problems, we employ data augmentation techniques such as random rotating by integer angles ($90^{\circ}$, $180^{\circ}$, $270^{\circ}$) and random horizontal/vertical flipping. These augmentations are applied to implicitly increase the data volume without altering the essence of Bongard-Logo as a few-shot learning task or explicitly changing the level of reasoning difficulty. We anticipate that this augmentation approach will enable PMoC and (Straw) Pose-Transformer to unleash their greater potential. The results is illustrated in Table II.
TABLE II: Reasoning Accuracies of PMoC on Augmented Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
PMoC 94.5 92.6 98.0 78.3 76.5
PMoC+Straw-Pose-Transformer 98.7 96.3 98.3 79.5 76.6
PMoC+Pose-Transformer Deployment is not feasible on our device
This paper argues that the adopted data augmentation methods do not explicitly alter the core reasoning difficulty or logical complexity of the Bongard-Logo problem. These augmentation techniques, from an intuitive perspective, neither add nor remove any preexisting concepts. Even if these data augmentation methods implicitly introduce additional and unknown concepts to the Bongard-Logo dataset, it is important to note that such concepts are not present in the test set, as no data augmentation has been applied to it. Furthermore, they do not provide guidance for generalization issues such as NV and CM, as the solutions to these problems are not directly reflected in the HD problem after applying such augmentations. Additionally, the data augmentation approach employed in this study is uniformly applied to all 14 images within each Bongard-Logo instance, maintaining the nature of few-shot learning (i.e., without increasing the number of images within an instance) and without introducing additional priors (such as rotational and flipping invariance of individual images within Bongard-Logo instances). These augmentation methods indeed enrich the expressions of concepts at the pixel level, which is precisely the challenge we anticipate for PMoC through data augmentation and expanded data volume. Enabling multiple expressions for concepts is precisely the original intention behind the establishment of the Bongard-Logo dataset[16].
VI-B Experiment on RPM
Given that I-RAVEN[19] represents a more advanced and rigorous benchmark compared to RAVEN[17], the experiments in this paper were solely conducted on the I-RAVEN dataset. We conducted experiments on I-RAVEN and PGM [18] by sequentially replacing the Transformer-Encoder structure within RS-Tran [30] with Pose-Transformer and Straw-Pose-Transformer to obtain RS-Pose-Tran and RS-Straw-Pose-Tran. The results of I-RAVEN are recorded in Table III and PGM is recorded in Table IV. In RS-Tran, multiple layers of the standard Transformer-Encoder are stacked. However, considering our belief in the significant performance advantages of (Straw) Pose-Transformer compared to the standard Transformer-Encoder, we only stacked half of the logical layers compared to RS-Tran when creating the combined models RS-Pose-Tran and RS-Straw-Pose-Tran. Additionally, the parameter settings and dataset configurations for experiments related to RS-Pose-Tran were maintained strictly consistent with those of RS-Tran. This setup ensures that the comparison results among the three models are more convincing.
TABLE III: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
SAVIR-T [26] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
SCL [25, 26] 91.6/95.0 98.1/99.0 91.0/96.2 82.5/89.5 96.8/97.9 96.5/97.1 96.0/97.6 80.1/87.7
MRNet [20] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[30] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
RS-Pose-TRAN -/99.51 -/100.00 -/99.90 -/98.20 -/100.00 -/100.00 -/99.94 -/98.56
RS-Straw-Pose-TRAN -/99.44 -/100.0 -/99.81 -/98.12 -/100.00 -/100.00 -/99.97 -/98.22
In Table III and IV, we observed that compared to the Transformer encoder, the Pose-Transformer exhibited a greater aptitude for abstract reasoning problems, particularly in addressing RPM and Bongard-Logo challenges. Furthermore, it is noteworthy that we refrained from conducting experiments with RS-Pose-Tran on the PGM dataset. This decision was made due to the fact that, compared to RS-Tran operating on the RAVEN dataset, the computational overhead of RS-Tran doubled when applied to the PGM dataset. However, such a significant increase in computational requirements was prohibitive for RS-Pose-Tran, rendering it unfeasible for deployment on our current equipment and facilities. Consequently, this motivated us to introduce the Straw-Pose-Transformer in this study, a lightweight variant of the Pose-Transformer.
TABLE IV: Reasoning Accuracies on PGM.
Model Test Accuracy(%)
SAVIR-T [26] 91.2
SCL [25, 26] 88.9
MRNet [20] 94.5
RS-CNN[30] 82.8
RS-TRAN[30] 97.5
RS-Straw-Pose-TRAN 98.6(98.64)
RS-Pose-TRAN Deployment is not feasible on our device
From the perspective of experimental results, it is evident that our lightweight approach, RS-Straw-Pose-Tran, has achieved success without compromising the performance of RS-Pose-Tran to a significant degree. Moreover, the incorporation of the pose mapping matrix into the Transformer-Encoder has proven effective, leading to a substantial enhancement in the performance of RS-Tran. Even without taking into account the improvements relative to RS-Tran, the results achieved by RS-Straw-Pose-Tran are indeed noteworthy and can be classified as remarkable.
VII Conclusion
The characteristics of deep networks as black-box intelligent systems, coupled with their optimization methods based on steepest gradient descent, create a gap between the theory and implementation of deep networks. In this paper, we report the design process of PMoC, which represents a “compromise” and “indirect” approach to neural network architecture design philosophy. By employing a unique probabilistic modeling technique, PMoC cleverly circumvents many of the theoretical and practical discrepancies associated with deep learning. This allows PMoC to offer a probabilistic-model-based solution for visual clustering problems. Its effectiveness has been validated on the renowned clustering reasoning problem: Bongard-Logo. This paper advocates for the adoption of similar indirect strategies and compromise approaches when addressing similar visual clustering problems or tasks requiring the estimation of unknown distributions.
As a bonus, we have designed the Pose-Transformer, a variant of the attention-based Transformer, which has excelled in logical reasoning tasks. Both RS-Tran[30] and the PMoC utilize the Transformer as their core backbone structure, and they have achieved notable successes in abstract reasoning problems. This demonstrates the remarkable logical analysis capability inherent in the Transformer. By replacing the Transformer structure in them with our Pose-Transformer, we have observed a direct and significant improvement in their accuracy. This paper encourages the consideration of our versatile hot-swappable structure when addressing visual abstract reasoning problems.